Grade 12Physical Chemistry

What is the value of vant hoff factor for a compound which goes tetramerisation in solvent

Manjari
7 Years agoGrade 12

2 Answers

Pooja
7 Years ago
van`t hoff factor = Number of moles of particle after association / Number of moles of particles before association
 
Here i=1/4

X??1/4X4

 

 
 
 
sree lakshmi
7 Years ago
  i = no.  Of moles after association /no.  Of moles                                                                           before association
In tetramerisation
4A\rightarrow A^{_{ _{4}}}
Ie     I=1/4 = 0.25